Home‎ > ‎Crew 133‎ > ‎

113-GreenHabReport

XO Gartrelle to Cap Com

The Greenhab is operating normally.

The fan is operating normally set at 68 degrees F.

The heater was set to 40 degrees F at 9AM
The heater was set to 75 degrees F at 5PM

Only the hydroponic seedlings received water today.The onion, basil and cilantro seedlings have not germinated. The Mars Shoot is fine. The hydroponic seedlings were placed in the green hab at 9AM and returned to the upper Main hab for the night at 5PM. 

The iniitial hydroponic procedures are completed. Commander Crock, Pedro Diaz-Rubin, Joe Jessup and Elizabeth Howell all assisted in the process. The procedure started at 2:50 and completed at 3:20.  We have placed 16 gallons of water in the reservoir, added chemicals, started the pump, positioned the overfill bucket and the meter is running (Meter was started at approximately 3:10). Syringes were rinsed and returned to their storage position. The chemical/water mixture will be agitated throughout the night. The lettuce shoots will be planted tomorrow morning. 

Photos from taken in the Greenhab today are attached. 

Greenhab readings for today are as follows:

                               Current              Max               Min   
Temperature         81 Degrees       99 Degrees         48 Degrees
Humidity              14%                 19 %                  10 %